# Gender Bias and Property Taxes

### v1

@@ 6.2 Descriptive Results
We now turn to Gemini’s responses about the appellants’ and panelists’ behaviors and tone. In Figure 4 we graphically depict the incidence rate for each behavior or tone for appellants and panelists, reporting split sample means and 95% confidence intervals for each combination of attendee gender and ARB chair gender. The values we depict are re-centered around the population mean incidence for the respective behavior or tone, to facilitate easier comparisons of cross-gender differences. For example, we observe that, in the case of male appellants facing male ARB chairs, appellants are systematically more likely to sound confident (they are labeled as such 1.75% more often than the average incidence across our entire sample). In contrast, female appellants are systematically more likely to ask questions of the panel when facing a female ARB chair, approximately 1.6% more often than the overall average.
@@
For the first analysis, we conduct four subsample regressions, splitting on combinations of appellant gender and ARB chair gender. When doing so, we focus on the AnySaved outcome, for the sake of simplicity. We also focus strictly on cases where Gemini’s gender predictions align with our regular-expression-based gender predictions to minimize potential measurement error. We work with approximately three-quarters of the audio-recorded sample (roughly 62,000 records across the four sample splits). We report the result of these four estimations below, in Table 14. We observe, in most cases, that appellants’ behaviors and tone have little to no impact on hearing outcomes, with a handful of exceptions. Note that we report these regressions in Table 14 to explore correlations; that is, the results do not necessarily represent causal effects. For example, we see that presenting formal evidence has a significant, positive relationship with home value reductions, and this effect is relatively homogeneous across all gender-pair sub-samples, increasing the probability of a value reduction by approximately 6-7 pp. Further, we observe that ’sounding confident’ has a positive impact on hearing outcomes specifically when ARB panels are chaired by men. We also observe that ’sounding assertive’ has a significant positive impact on hearing outcomes, specifically for male appellants. Importantly, however, the only effect we see that would suggest an explanation for why females fare worse (better) in front of female (male) ARB chairs is that female appellants obtain a significant improvement in their outcomes when they ’sound defensive’ in front of a female ARB chair, yet not when they face a male ARB chair. A particularly salient takeaway from these estimates is that, despite our observation that males are systematically more aggressive in their behavior in front of female ARB chairs (raising their voice, interrupting, disagreeing, sounding hostile), this has no material consequence for their hearing outcomes. This suggests that the behavior is tolerated and perhaps even expected.
@@
App Raises Voice 0.009 (0.031) -0.007 (0.017) 0.016 (0.029) 0.021 (0.018)
App Disagrees -0.014 (0.013) -0.030*** (0.008) -0.018 (0.016) -0.021* (0.011)
App Interrupts -0.009 (0.012) 0.012 (0.008) -0.006 (0.015) 0.004 (0.009)
App Refers to Evidence 0.062*** (0.010) 0.063*** (0.007) 0.065*** (0.012) 0.069*** (0.008)
App Sounds Confident 0.012 (0.017) 0.009 (0.012) 0.041*** (0.019) 0.026*** (0.012)
App Sounds Frustrated 0.007 (0.018) -0.014 (0.009) 0.027 (0.019) -0.012 (0.011)
App Sounds Defensive 0.023* (0.013) -0.010 (0.009) -0.003 (0.013) -0.010 (0.010)
App Sounds Assertive 0.013 (0.024) 0.023* (0.013) -0.001 (0.028) 0.049*** (0.016)
App Sounds Nervous 0.001 (0.028) 0.015 (0.024) 0.033 (0.026) 0.065*** (0.030)
App Sounds Angry 0.022 (0.043) 0.015 (0.026) 0.017 (0.038) -0.031 (0.028)
App Sounds Accusatory -0.048 (0.041) -0.027 (0.021) -0.027 (0.041) 0.000 (0.028)
App Asks Questions -0.011 (0.012) 0.010 (0.008) 0.003 (0.013) 0.008 (0.008)
App Asks Repeat -0.004 (0.012) 0.003 (0.007) -0.005 (0.013) -0.011 (0.009)
App Asks Speak Louder -0.003 (0.023) 0.025* (0.013) -0.017 (0.024) -0.007 (0.017)
@@
App Refers to Evidence 0.067*** (0.008) 0.069*** (0.006) 0.062*** (0.008) 0.064*** (0.006)
App Sounds Confident 0.027** (0.012) 0.016** (0.008) 0.020 (0.013) 0.011 (0.008)
App Sounds Frustrated 0.022* (0.013) -0.009 (0.007) 0.024* (0.013) -0.006 (0.007)
App Sounds Defensive 0.009 (0.009) -0.008 (0.007) 0.005 (0.009) -0.012* (0.007)
App Sounds Assertive -0.001 (0.017) 0.036*** (0.009) -0.003 (0.018) 0.036*** (0.010)
App Sounds Nervous 0.006 (0.019) 0.027 (0.019) -0.005 (0.019) 0.029 (0.019)
App Sounds Angry 0.025 (0.029) -0.002 (0.019) 0.022 (0.029) -0.006 (0.019)
App Sounds Accusatory -0.034 (0.030) -0.018 (0.017) -0.038 (0.030) -0.014 (0.017)
App Asks Questions -0.004 (0.009) 0.008 (0.006) 0.006 (0.009) 0.011* (0.006)
App Asks Repeat -0.004 (0.009) -0.006 (0.006) -0.004 (0.009) -0.005 (0.006)
App Asks Speak Louder -0.004 (0.016) 0.012 (0.011) -0.007 (0.017) 0.011 (0.011)
@@
Female Panel Chair -0.168** (0.076) -0.018 (0.051)
Female Panelists (1) -0.250** (0.110) 0.013 (0.083)
@@
Female Panelists (3) -0.388*** (0.127) -0.005 (0.092)
Market Equity 1.427*** (0.183) 1.482*** (0.101) 1.448*** (0.183) 1.486*** (0.101)
Unequal Appraisal 0.195** (0.080) 0.224*** (0.056) 0.213*** (0.080) 0.222*** (0.056)
Online -0.443*** (0.140) -0.202* (0.109) -0.396** (0.166) -0.131 (0.117)
@@
Saved Last Protest 0.752*** (0.100) 0.693*** (0.067) 0.711*** (0.103) 0.715*** (0.065)
@@
App Raises Voice 0.062 (0.196) 0.069 (0.129) 0.149 (0.203) 0.067 (0.129)
App Disagrees -0.165* (0.099) -0.318*** (0.068) -0.142 (0.107) -0.288*** (0.074)
App Interrupts -0.130 (0.093) 0.057 (0.062) -0.124 (0.095) 0.052 (0.064)
App Refers to Evidence 0.856*** (0.074) 0.856*** (0.054) 0.813*** (0.078) 0.812*** (0.053)
App Sounds Confident 0.252** (0.116) 0.195** (0.077) 0.183 (0.126) 0.146* (0.078)
App Sounds Frustrated 0.329*** (0.123) 0.025 (0.070) 0.347*** (0.126) 0.056 (0.073)
App Sounds Defensive 0.108 (0.088) -0.067 (0.066) 0.073 (0.091) -0.110 (0.067)
App Sounds Assertive 0.019 (0.167) 0.368*** (0.091) 0.011 (0.175) 0.369*** (0.097)
App Sounds Nervous 0.143 (0.190) 0.259 (0.184) 0.011 (0.190) 0.256 (0.184)
App Sounds Angry 0.215 (0.286) 0.063 (0.190) 0.194 (0.286) 0.045 (0.193)
App Sounds Accusatory -0.344 (0.297) -0.270 (0.171) -0.387 (0.292) -0.239 (0.173)
App Asks Questions -0.011 (0.086) 0.155** (0.062) 0.076 (0.091) 0.188*** (0.064)
App Asks Repeat 0.007 (0.087) -0.026 (0.055) 0.006 (0.092) -0.011 (0.056)
App Asks Speak Louder -0.068 (0.160) 0.084 (0.102) -0.102 (0.168) 0.070 (0.107)
@@
Finally, we repeat the estimation accounting for high-dimensional embedded representations of the broader, unstructured content of hearing audio recordings, rather than conditioning upon specific Gemini-coded features. We achieve this employing a double machine learning framework (Chernozhukov et al., 2018), an approach that has been taken in other recent literature to account for unstructured data in a relatively non-parametric manner (Manzoor et al., 2024). The double machine learning framework is useful here because it simultaneously enables us to accommodate many covariates, i.e., our embedding dimensions and their likely non-linear interactive relationships with appellant gender, panelist gender, and hearing outcomes. Our implementation models the relationships between our nuisance parameters (structured covariates and text-embedding dimensions), the outcome, and the ARB panelist gender via a pair of random forest models. The text-embedding for each hearing audio recording was obtained by first passing the recording through Open AI’s Whisper Medium speech-to-text model. Because Whisper model inference is relatively slow (compared to inference with Gemini Flash, at least), we conduct this exercise focusing strictly on hearing audio records from 2022. The textual transcription for each hearing recording was then passed through Google’s text-embedding-004 model, yielding a high-dimensional vector representation of the spoken content. Finally, each embedding dimension is treated as a distinct covariate in our estimation, which we implement employing the DoubleML package in R, specifically its implementation of a partial linear specification. We focus on the linear probability model that considers the Any Saved outcome. The results of these estimations are presented in Table 17, where we again observe consistent results.
@@ 7 Discussion
Our work is, of course, subject to several limitations. First, we focus on a single geography, Harris, County, Texas. As such, generalizability is a question. Houston may have a unique sociocultural climate, and the local market may be subject to idiosyncrasies in terms of norms and laws, as well as its property appraisal appeals process. That said we have observed broad evidence of gender-concordant bias across the entire pool of ARB panelists (see Figure 1), suggesting this phenomenon may be more general, cutting across political and social groups. We can evaluate the generalizability of our results using the SANS framework (selection, attrition, naturalness, and scalability) developed in (List, 2020). Regarding selection, our sample is representative of hearings in Harris County, the target population. However, we note potential selection bias, as females may be less likely to participate in hearings due to anticipated biases against them. Notably, in our setting, participants do not know the gender of the panel members before the hearing. Given the similarity of hearings across Texas’s 241 counties, our sample may also be representative of hearings statewide. Considering attrition, experiences, and outcomes in prior hearings could influence individuals’ willingness to participate in future hearings, potentially affecting the composition of future samples. Regarding naturalness, both homeowners and panelists operate within a natural margin, engaging in real decisions rather than an artificial task designed for research purposes. Regarding scalability, there are no clear reasons to anticipate a voltage drop if the analysis is scaled to other counties in Texas. However, once our findings on biases in this context become public, behavior may change. For instance, the Texas state comptroller, who oversees hearings in all counties in the state, may share our findings with county ARBs, and appellants might adjust their behavior in future hearings. In sum, we view our findings as a WAVE1 insight in the nomenclature of (List, 2020). Replications are necessary to determine whether our results generalize to other administrative processes, and future work can and should explore these dynamics by leveraging administrative data from other areas of the United States.
@@
Third, while we leverage a state-of-the-art multi-modal large language model (Gemini 1.5 Flash) to analyze audio recordings, the model’s inferences about tone and behavior are not without error. Although we validate some of these inferences against known ground truths, there remains the possibility that certain nuances of human interaction—especially those influenced by socio-cultural context—are missed or misinterpreted by the model. This limitation suggests that human coding, or a hybrid approach combining machine learning with human validation may be a valuable avenue for future research.
@@ 8 Conclusion
Beyond these contributions, our results could also affect future behavior by appellants and panelists, reducing biases, in which case they would represent a public service. Indeed, when individuals know they are being recorded, their behavior often changes due to the observer effect, where heightened scrutiny influences actions. Teachers, for instance, tend to modify their teaching practices during evaluations, focusing more on structured lessons and student engagement to meet professional standards (Kane et al., 2013). Similarly, judges may exhibit greater formality in court proceedings to align with public and appellate expectations, especially when media coverage is present (Lim, Snyder and Strömberg, 2015). Legislators similarly adjust their rhetoric during televised sessions, emphasizing performative or emotional appeals to resonate with constituents (Osnabrugge, Hobolt and Rodon, 2021). This shift toward transparency and visibility creates both opportunities for accountability and challenges in maintaining authentic behavior. Research consistently shows that such surveillance amplifies the emphasis on appearance and messaging, often at the expense of substance (Ash and MacLeod, 2017; Boussalis et al., 2021). Finally, our paper illustrates how advancements in AI can enable the analysis of recordings and data previously deemed unsuitable for examination, uncovering insights akin to how DNA revolutionized cold case investigations and helped to solve decades-old mysteries.
@@
For each audio recording, we obtain annotations using Google’s ’genai’ API (Application Programming Interface), using Python. All API requests employed ’gemini-1.5-flash-001’ and requested a JSON formatted response. Audio annotations for hearings from 2022 were obtained in late August of 2024. Additional hearing records were obtained for 2013 through 2021 in early October of 2024 and were coded in the latter half of that month. For every recording, a zero-shot prompt was employed, using the default temperature setting (a value of 1.0). The exact prompt for the coding exercise was as follows.
@@
  "board_member_who_speaks_most_gender": "ANSWER MALE or FEMALE",
  "count_people_present_for_property_owner": "ANSWER WITH A NUMBER, e.g., 2",
@@
  "appellant_ever_sounds_accusatory_or_hostile": "ANSWER YES or NO",
@@
  "appellant_type_of_evidence_if_any":
"ANSWER WITH FREE FORM TEXT OR NOT APPLICABLE",
  "appellant_property_owner_or_hired_agent":
"ANSWER WITH OWNER or AGENT",
  "appraiser_gender": "ANSWER WITH MALE or FEMALE",
  "appellant_time_speaking_during_hearing":
  "ANSWER WITH A PERCENTAGE VALUE, e.g., 50%",

### v2

@@ 6.1 The Quality of LLMs’ Audio Coding
Note that we also obtained annotations from human coders for a random sample of recordings, for comparison. That analysis, reported in Appendix B, suggests that Gemini’s annotations do contain meaningful information.
@@ 6.2 Descriptive Results
We now turn to Gemini’s responses about the appellants’ and panelists’ behaviors and tone. In Figure 4 we graphically depict the incidence rate for each behavior or tone for appellants and panelists, reporting split sample means and 95% confidence intervals for each combination of attendee gender and ARB chair gender. The values we depict are re-centered around the population mean incidence for the respective behavior or tone, to facilitate easier comparisons of cross-gender differences. For example, we observe that, in the case of male appellants facing male ARB chairs, appellants are systematically more likely to sound confident (they are labeled as such 1.75% more often than the average incidence across our entire sample). In contrast, female appellants are systematically less likely to sound confident when facing a male ARB chair. This annotation occurs nearly 2% less often than the overall incidence in the population.
@@
For the first analysis, we conduct four subsample regressions, splitting on combinations of appellant gender and ARB chair gender, to explore correlations; that is, the estimates from these regressions do not necessarily represent causal effects. We focus on the AnySaved outcome in these analyses, for the sake of simplicity. We also focus strictly on cases where Gemini’s gender predictions align with our regular-expression-based gender predictions to minimize potential measurement error. Thus, we work with approximately three-quarters of the audio-recorded sample (roughly 62,000 records across the four sample splits). Our results are shown in Table 14, where we see a few interesting associations. For example, we see that presenting formal evidence has a significant, positive relationship with home value reductions, and this effect is relatively homogeneous across all gender-pair sub-samples, increasing the probability of a value reduction by approximately 6-7 pp. Further, we observe that ’sounding confident’ has a positive impact on hearing outcomes specifically when ARB panels are chaired by men, and that male appellants, in particular, are penalized for disagreeing verbally with the panel. We also observe that ‘sounding hostile’ has a significant negative impact on hearing outcomes, specifically when ARB panels are chaired by women. Importantly, however, we see no effects that would suggest an explanation for why females fare worse (better) in front of female (male) ARB chairs. A particularly salient takeaway from these estimates is that, despite our observation that males are systematically more aggressive in their behavior in front of female ARB chairs (raising their voice, interrupting, disagreeing, sounding hostile), this has very limited material consequences for their hearing outcomes. This suggests that the behavior is tolerated and perhaps even expected.
@@
App Raises Voice 0.016 (0.032) -0.005 (0.017) 0.021 (0.029) 0.033* (0.020)
App Disagrees -0.015 (0.014) -0.029*** (0.009) -0.023 (0.015) -0.022** (0.011)
App Interrupts -0.010 (0.013) 0.012 (0.008) -0.007 (0.003) 0.004 (0.009)
App Refers to Evidence 0.069*** (0.011) 0.067*** (0.007) 0.073*** (0.012) 0.070*** (0.008)
App Sounds Confident 0.013 (0.017) 0.012 (0.012) 0.040** (0.017) 0.028** (0.012)
App Sounds Frustrated 0.018 (0.018) -0.016* (0.010) 0.012 (0.017) -0.010 (0.011)
App Sounds Defensive 0.036*** (0.013) -0.001 (0.009) 0.007 (0.014) -0.001 (0.011)
App Sounds Nervous -0.009 (0.027) 0.003 (0.025) 0.025 (0.027) 0.057* (0.030)
App Sounds Angry 0.041 (0.042) 0.023 (0.026) 0.014 (0.039) -0.024 (0.031)
App Sounds Hostile -0.071* (0.040) -0.041* (0.021) -0.022 (0.040) -0.015 (0.030)
App Asks Questions -0.006 (0.012) 0.015* (0.008) 0.003 (0.014) 0.007 (0.008)
App Asks Repeat -0.009 (0.012) -0.002 (0.007) -0.009 (0.013) -0.016* (0.009)
App Asks Speak Louder -0.005 (0.024) 0.020 (0.014) -0.022 (0.025) -0.010 (0.017)
@@
App Refers to Evidence 0.067*** (0.008) 0.071*** (0.006) 0.062*** (0.008) 0.066*** (0.006)
App Sounds Confident 0.027** (0.012) 0.021** (0.008) 0.020 (0.013) 0.016** (0.008)
App Sounds Frustrated 0.022* (0.013) -0.004 (0.007) 0.024* (0.013) -0.001 (0.007)
App Sounds Defensive 0.009 (0.009) -0.007 (0.007) 0.005 (0.009) -0.011 (0.007)
App Sounds Nervous 0.006 (0.019) 0.027 (0.019) -0.005 (0.019) 0.028 (0.019)
App Sounds Angry 0.025 (0.029) -0.002 (0.019) 0.021 (0.029) -0.006 (0.019)
App Sounds Hostile -0.034 (0.030) -0.019 (0.017) -0.038 (0.030) -0.014 (0.017)
App Asks Questions -0.004 (0.009) 0.009 (0.006) 0.006 (0.009) 0.012* (0.006)
App Asks Repeat -0.004 (0.009) -0.005 (0.006) -0.004 (0.009) -0.005 (0.006)
App Asks Speak Louder -0.004 (0.016) 0.011 (0.011) -0.007 (0.017) 0.011 (0.011)
@@
Female Panel Chair -0.169** (0.076) -0.018 (0.051)
Female Panelists (1) -0.250** (0.110) 0.011 (0.083)
@@
Female Panelists (3) -0.388*** (0.127) -0.006 (0.092)
Market Equity 1.426*** (0.182) 1.476*** (0.101) 1.448*** (0.182) 1.481*** (0.101)
Unequal Appraisal 0.194** (0.080) 0.220*** (0.056) 0.213*** (0.080) 0.218*** (0.056)
Online -0.443*** (0.140) -0.199* (0.108) -0.396** (0.166) -0.132 (0.117)
@@
Saved Last Protest 0.752*** (0.100) 0.694*** (0.067) 0.711*** (0.103) 0.716*** (0.065)
@@
App Raises Voice 0.062 (0.196) 0.071 (0.129) 0.149 (0.203) 0.070 (0.129)
App Disagrees -0.164* (0.097) -0.276*** (0.067) -0.141 (0.105) -0.245*** (0.074)
App Interrupts -0.131 (0.093) 0.054 (0.062) -0.124 (0.095) 0.048 (0.064)
App Refers to Evidence 0.856*** (0.074) 0.880*** (0.054) 0.814*** (0.078) 0.836*** (0.053)
App Sounds Confident 0.255** (0.118) 0.247** (0.076) 0.184 (0.127) 0.199** (0.077)
App Sounds Frustrated 0.332*** (0.119) 0.074 (0.070) 0.349*** (0.123) 0.105 (0.072)
App Sounds Defensive 0.109 (0.089) -0.060 (0.066) 0.073 (0.091) -0.103 (0.067)
App Sounds Nervous 0.144 (0.191) 0.254 (0.183) 0.011 (0.190) 0.252 (0.183)
App Sounds Angry 0.215 (0.286) 0.069 (0.190) 0.195 (0.286) 0.052 (0.193)
App Sounds Hostile -0.344 (0.297) -0.277 (0.171) -0.387 (0.292) -0.247 (0.173)
App Asks Questions -0.011 (0.086) 0.169*** (0.062) 0.077 (0.092) 0.202*** (0.063)
App Asks Repeat 0.007 (0.087) -0.020 (0.055) 0.006 (0.092) -0.004 (0.056)
App Asks Speak Louder -0.068 (0.160) 0.080 (0.102) -0.102 (0.168) 0.066 (0.107)
@@
Finally, we repeat the estimation accounting for high-dimensional embedded representations of the broader, unstructured content of hearing audio recordings, rather than conditioning upon specific Gemini-coded features. We achieve this employing a double machine learning framework (Chernozhukov et al., 2018), an approach that has been taken in other recent literature to account for unstructured data in a relatively data-driven, non-parametric manner (Manzoor et al., 2024). The double machine learning framework is useful here because it simultaneously enables us to accommodate many covariates, i.e., our embedding dimensions, and their likely non-linear interactive relationships with appellant gender, panelist gender, and hearing outcomes. We modeled the relationships between our nuisance parameters (structured covariates and text-embedding dimensions), the outcome, and the ARB panelist gender via a pair of random forest models. The text-embedding for each hearing audio recording was obtained by first passing the recording through Open AI’s Whisper Medium speech-to-text model. Because Whisper model inference is relatively slow (compared to inference with Gemini Flash, at least), we conduct this exercise focusing strictly on hearing audio records from 2022. The textual transcription for each hearing recording was then passed through Google’s ‘text-embedding-004’ model, yielding a 768-dimensional vector representation of the spoken content. [^17] Stack Overflow Blog — “An Intuitive Introduction to Text Embeddings”: https://stackoverflow.blog/2023/11/09/an-intuitive-introduction-to-text-embeddings/ Finally, each embedding dimension is treated as a distinct covariate in our estimation, which we implement employing the DoubleML package in R, specifically its implementation of a partial linear specification. We focus on the linear probability model that considers the Any Saved outcome. The results of these estimations are presented in Table 17, where we again observe consistent results.
@@ 7 Discussion
Our work is, of course, subject to several limitations. First, we focus on a single geography, Harris, County, Texas. As such, generalizability is a question. Houston may have a unique sociocultural climate, and the local market may be subject to idiosyncrasies in terms of norms and laws, as well as its property appraisal appeals process. That said we have observed broad evidence of gender-concordant bias across the entire pool of ARB panelists (see Figure 1), suggesting this phenomenon may be more general, cutting across political and social groups. We can evaluate the generalizability of our results using the SANS framework (selection, attrition, naturalness, and scalability) developed in (List, 2020). Regarding selection, our sample is representative of hearings conducted without the involvement of professional agents in Harris County, the target population. However, we note potential selection bias, as females may be less likely to participate in hearings due to anticipated biases against them. Notably, in our setting, participants do not know the gender of the panel members before the hearing. Given the similarity of hearings across Texas’s 241 counties, our sample may also be broadly representative of hearings conducted without professional agents across the state. Considering attrition, experiences, and outcomes in prior hearings could influence individuals’ willingness to participate in future hearings, potentially affecting the composition of future samples. Regarding naturalness, both homeowners and panelists operate within a natural margin, engaging in real decisions rather than an artificial task designed for research purposes. Regarding scalability, there are no clear reasons to anticipate a voltage drop if the analysis is scaled to other counties in Texas. However, once our findings on biases in this context become public, behavior may change.
In sum, we view our findings as a WAVE1 insight in the nomenclature of (List, 2020). Replications are necessary to determine whether our results generalize to other administrative processes, and future work can and should explore these dynamics by leveraging administrative data from other areas of the United States.
@@
Third, while we leverage a state-of-the-art multi-modal large language model (Gemini 1.5 Flash) to analyze audio recordings, the model’s inferences about tone and behavior are not without error. Although we validate some of these inferences against known ground truths from administrative data, and we also compared these inferences with annotations from human coders for a random sample of recordings (see Appendix B), there remains the possibility that certain nuances of human interaction—especially those influenced by socio-cultural context—are missed or misinterpreted by the model. This limitation suggests that human coding, or a hybrid approach combining machine learning with human validation may be a valuable avenue for future research.
@@ 8 Conclusion
Beyond these contributions, our results could also affect future behavior by appellants and panelists, reducing biases, in which case they would represent a public service. Indeed, when individuals know they are being recorded, their behavior often changes due to the observer effect, where heightened scrutiny influences actions. Teachers, for instance, tend to modify their teaching practices during evaluations, focusing more on structured lessons and student engagement to meet professional standards (Kane et al., 2013). Similarly, judges may exhibit greater formality in court proceedings to align with public and appellate expectations, especially when media coverage is present (Lim, Snyder and Strömberg, 2015). Legislators similarly adjust their rhetoric during televised sessions, emphasizing performative or emotional appeals to resonate with constituents (Osnabrugge, Hobolt and Rodon, 2021). This shift toward transparency and visibility creates both opportunities for accountability and challenges in maintaining authentic behavior. Research consistently shows that such surveillance amplifies the emphasis on appearance and messaging, often at the expense of substance (Ash and MacLeod, 2017; Boussalis et al., 2021). Furthermore, our paper illustrates how advancements in AI can enable the analysis of recordings and data previously deemed unsuitable for examination, uncovering insights akin to how DNA revolutionized cold case investigations and helped to solve decades-old mysteries.
Finally, the dissemination of our results could have both immediate effects on property tax protest hearings in Texas and inspire a redesign of the administrative process in the future. Regarding immediate effects, the county appraisal districts in Texas (recall there are 241 counties), the Texas comptroller overseeing these county appraisal districts, or the panelists themselves could adjust their procedures and behavior to mitigate the biases we document. As for redesigning the administrative process, hearings could incorporate AI consulting bots to prevent biases and assist in adjudication decisions. Indeed, in follow-up work, we are exploring how AI can enhance the adjudication process.
@@
For each audio recording, we obtain annotations using Google’s ‘genai’ API (Application Programming Interface), using Python. All API requests employed ‘gemini-1.5-flash-001’ and requested a JSON formatted response. Audio annotations for hearings from 2022 were obtained in late August of 2024. Additional hearing records were obtained for 2013 through 2021 in early October of 2024 and were coded in the latter half of that month. For every recording, a zero-shot prompt was employed, using the default temperature setting (a value of 1.0). The exact prompt for the coding exercise was as follows.
@@
  "appellant_property_owner_or_hired_agent": "ANSWER WITH OWNER or AGENT",
@@
  "appellant_ever_sounds_accusatory_or_hostile": ANSWER YES or NO",